# On the Gravitational Wave Counterpart to a Gamma-ray Galactic Center Signal from Millisecond Pulsars

### v1

@@ I Introduction
Unlike DM annihilation, MSPs are slated to produce a gravitational wave signal; the latter originates from the non-spherical nature of the rotating neutron star, a consequence, in turn, of its large rotational speeds (see e.g. Agarwal et al. (2022) and references therein). Thus, detecting the gravitational wave (GW) counterpart generated by MSPs could potentially solve the open question of the origin of the GCE, as pointed out in a few previous analyses Calore et al. (2019); Agarwal et al. (2022); Miller and Zhao (2023). For instance, in the present study, we re-assess the connection between the GCE and GW emission in the MSPs scenario for the GCE, utilizing state-of-the-art population models and statistical methods.
@@ II Millisecond Pulsar Models
$\epsilon\approx 10^{-8}\left(\frac{B_{int}}{10^{12}}\right).$ (2)
@@
with $d$ the distance to the MSP from Earth, $f_{rot}$ the MSP’s rotational frequency, and $G$ the gravitational constant.
@@ III.1 Millisecond Pulsar Densities in the Milky Way
The Boxy model includes two sub-populations: a disk and a nuclear bulge population. The disk population, assumed to only depend on the Galactic latitude $z$ and radius $R$ coordinates, is assumed to follow a density of the form:
@@
Figure 1: A colormap of the MSP “Boxy” density distribution in the Milky Way disk, in log10 scale; Left: the disk component; Right: the nuclear bulge component.
The nuclear bulge distribution is, in turn, split into two separate populations, that of the nuclear stellar cluster (NSC) and that of the nuclear stellar disk (NSD):
$\rho_{\text{nuclear bulge}}(r,z)=\rho_{\text{NSC}}(r)+\rho_{\text{NSD}}(r,z)$ (6)
with $r^{2}=x^{2}+y^{2}+z^{2}$ Ploeg (2021). The nuclear stellar cluster is assumed to have a density:
$\rho_{\text{NSC}}(r)=\begin{cases}\frac{\rho_{\text{0, NSC}}}{1+\left(\frac{r}{r_{0}}\right)^{2}}&r\leq 6\text{ pc}\\ \frac{\rho_{\text{1, NSC}}}{1+\left(\frac{r}{r_{0}}\right)^{3}}&6\text{ pc}<r\leq 200\text{ pc}\\ 0&r>200\text{ pc}\end{cases}$ (7)
with $r_{0}$ = 0.22 pc, $\rho_{\text{0, NSC}}$ = 3.3 $\times$ $10^{6}$ $M\textsubscript{$\odot$}\text{pc}^{-3}$, and $\rho_{\text{1, NSC}}$ = 8.9 $\times$ $10^{7}$ $M\textsubscript{$\odot$}\text{pc}^{-3}$ Ploeg (2021). The nuclear stellar disk density distribution is instead assumed to have the following functional form:
$\rho_{\text{NSD}}(r,z)=\begin{cases}\rho_{\text{0, NSD}}\left(\frac{r^{-0.1}}{1\text{ pc}}\right)\exp{\frac{-|z|}{45\text{ pc}}}&r<120\text{ pc}\\ \rho_{\text{1, NSD}}\left(\frac{r^{-3.5}}{1\text{ pc}}\right)\exp{\frac{-|z|}{45\text{ pc}}}&120\text{ pc}\leq r<220\text{ pc}\\ \rho_{\text{2, NSD}}\left(\frac{r^{-10}}{1\text{ pc}}\right)\exp{\frac{-|z|}{45\text{ pc}}}&r\geq 220\text{ pc}\end{cases}$ (8)
with $\rho_{\text{0, NSD}}$ = 301 $M\textsubscript{$\odot$}\text{pc}^{-3}$, $\rho_{\text{1, NSD}}$ = 3.53 $\times$ $10^{9}$ $M\textsubscript{$\odot$}\text{pc}^{-3}$, and $\rho_{\text{2, NSD}}$ = 5.94 $\times$ $10^{24}$ $M\textsubscript{$\odot$}\text{pc}^{-3}$ Ploeg (2021). See Fig. 1, right, for a plot of the nuclear bulge distribution.
The second, spherical population of MSPs consists of a single, spherically symmetric density distribution, given by:
@@
Inverse Transform Sampling is another way to sample from a certain distribution $p(x)$, provided its cumulative distribution function (CDF) is known. The CDF, evaluated at $y$, represents the probability that $p(x)$ will take a value less than or equal to y. Upon inversion, the inverse CDF, $I_{CDF}(x)=CDF^{-1}(x)$, can be used to generate a value from $p(x)$ when given an input in the range [0, 1] Hyvonen (2011).
Sampling a 2D function, f(x, y), requires splitting the function into two conditional functions, $f(x|y)$ and $f(y|x)$, as the MCMC and Inverse Transform Sampling methods discussed above are 1-dimensional. For piece-wise functions f(x,y), the integral bounds for $f_{x}(x)$ and $f_{y}(y)$ are given by the piece-wise bounds.
@@ III.2 Sampling the Millisecond Pulsar Population
with z, r, $z_{0},\sigma_{R}$ defined as in Eq. (5). A comparison between the generated sample, plotted as a 2D histogram, and Eq. (5), plotted as a contour map, can be seen in Fig. 3. To produce the spherical population, Eq. (9) was also sampled via MCMC methods, and was confirmed to match with the density distribution pictured in Fig. 2.
@@
The nuclear bulge populations were sampled using both MCMC and Inverse Transform methods as certain pieces of the nuclear bulge population, being rapidly decaying power laws, resulted in the MCMC algorithm ineffectively exploring the appropriate range of parameter space. For such pieces, Inverse Transform Sampling was used. The NSC density, consisting of the two equations given in Eq. (7), was sampled using MCMC methods. In addition to the NSC population, MCMC methods were used to sample parts of the NSD population. The following conditional distributions were sampled:
$\rho_{\text{NSD}}(z|r)\sim\exp{\left(\frac{-|z|}{45\ {\rm pc}}\right)}$ ) (12)
and
$\rho_{\text{NSD, 1}}(r|z)\sim\left(r/(1\ {\rm pc})\right)^{-0.1}$ (13)
with Eq. (13) in the range of [0, 120], and $z$, $r$ as defined in Eq. (8).
The last pieces of the NSD population were sampled using Inverse Transform methods. The general form of the inverse CDF for a power law of the form $f(x)=cx^{-a}$ given between two bounds $[x_{0},x_{1}]$ is:
$CDF^{-1}(x)=[(x_{1}^{1-a}-x_{0}^{1-a})y+x_{0}^{1-a}]^{\frac{1}{1-a}}$ (14)
with $a=3.5$, $x_{0}$ = 120 pc, and $x_{1}$ = 220 pc for the 2nd piece of the radial density in the NSD and $a=10$, $x_{0}$ = 220 pc, and $x_{1}$ = 500 pc for the 3rd piece of the radial density in the NSD. All sampled distributions were verified to match the functional forms described in Eq. (5), (7), and (8).
@@ IV Ellipticity and Moment of Inertia Distribution
The ellipticity distribution was then folded with the morphological models described above, resulting in the projected pulsar ellipticity distribution shown in Fig. 5 for a randomly drawn sample of 40,000 MSPs for the Boxy (left) and Spherical (right) population models. Note that the resultant ellipticity distribution is within the minimum ellipticity ranges described by Woan et al. Woan et al. (2018) and Chen et al. Chen (2020) of $\epsilon\approx 10^{-9}$, and agrees with the values obtained by Miller et al. Miller and Zhao (2023), as well as within the recent constraints given by Holst et al. Holst and Hooper (2024).
Figure 5: Locations and ellipticities of sampled MSPs for the boxy (left) and spherical (right) population models.
@@ VI Analysis and Possibility of Detection
Refer to captionRefer to caption
@@ VII Conclusion
In the future this analysis could be further strengthened by exploring the impact of a time-dependent ellipticity distribution and a frequency-dependent moment of inertia distribution on the strength and detectability of the GW signal. Additionally, it could be that the Galactic center excess is caused by a combination of MSPs and other, more exotic sources such as annihilating dark matter, or other unresolved astrophysical components, so it would be illuminating to incorporate such scenarios into the current model and distribution of MSPs and to evaluate the resulting GW signal. It should also be noted that Holst et al. Holst and Hooper (2024) suggests a range of possible MSP population sizes from 36,000-47,000 which, along with the unknown MSP luminosity function, suggests that the relation between the GCE and MSP population sizes is relatively poorly constrained. Therefore, further studies into how different population sizes affects detectability would be of great benefit.
Acknowledgements

### v2

@@ I Introduction
The predominant formation mechanism for MSPs in the Galactic center is believed to be associate with the so-called “cycling” formation scenario, in which a slowly rotating neutron star formed via a core-collapse supernova accretes material from a companion star during a low mass x-ray binary (LMXB) phase and is thus spun up into a MSP Lorimer (2008); Gautam et al. (2022). However, this formation mechanism has been found to predict too many LMXB systems in the Galactic center, casting doubt on its validity Gautam et al. (2022). Recent work has instead suggested that a population of MSPs resulting from the accretion-induced collapse (AIC) of O-Ne white dwarfs in the Galactic bulge is a significantly more likely pathway – both producing the correct intensity and morphology of the GCE signal while also remaining within LMXB constraints Gautam et al. (2022). Additional work also suggests that individual MSPs, found to be approximately 20-27% of MSPs in the Galactic center, are unlikely to contribute to the GCE signal; instead, binary systems are the likely culprit Macias et al. (2019); Jiang et al. (2020); Chanlaridis et al. (2024). The stellar encounter rate is significantly higher in the Galactic center than in the field, suggesting that the 20-27% range could potentially be an empirical upper limit McTier et al. (2020).
@@
Unlike DM annihilation, MSPs are slated to produce a gravitational wave signal; the latter originates from the non-spherical nature of the rotating neutron star, a consequence, in turn, of its large rotational speeds (see e.g. Agarwal et al. (2022) and references therein). Thus, detecting the gravitational wave (GW) counterpart generated by MSPs could potentially solve the open question of the origin of the GCE, as pointed out in a few previous analyses Calore et al. (2019); Agarwal et al. (2022); Miller and Zhao (2023); Carleo and Ben-Salem (2023). For instance, in the present study, we re-assess the connection between the GCE and GW emission in the MSPs scenario for the GCE, utilizing state-of-the-art population models and statistical methods.
@@ II Millisecond Pulsar Models
$\epsilon\approx 10^{-8}\left(\frac{B_{int}}{10^{12}{\rm\ G}}\right).$ (2)
@@
with $d$ the distance to the MSP from Earth, $f_{rot}$ the MSP’s rotational frequency, and $G$ the gravitational constant. These models assume that the ellipticities, and corresponding gravitational wave production, arise from isolated MSPs or MSPs with no orbital eccentricities–a simplifying assumption that we continue to adopt in the following work.
@@ III.1 Millisecond Pulsar Densities in the Milky Way
The Boxy model includes two sub-populations: a disk and a boxy bulge population, such that $\rho_{\text{boxy pop.}}=\rho_{\text{disk}}+\rho_{\text{boxy bulge}}$. The disk population, assumed to only depend on the Galactic latitude $z$ and radius $R$ coordinates, is assumed to follow a density of the form:
@@
Figure 1: A colormap of the "Boxy" MSP density distribution in the Milky Way disk, in log10 scale; Left: the disk component; Right: the boxy bulge component.
The boxy bulge distribution is, following Ploeg et al. (2020), given by:
$\rho_{\text{boxy bulge}}(R_{s})\propto\text{sech}^{2}(R_{s})\times\begin{cases}1&R\leq R_{\text{end}}\\ \exp{-\frac{(R-R_{\text{end}})^{2}}{h_{\text{end}}^{2}}}&R>R_{\text{end}}\\ \end{cases}$ (6)
with $R$ defined as above, $R_{\text{end}}=3128$ pc, and $h_{\text{end}}=461$ pc, and:
$R_{\perp}^{C_{\perp}}=\left(\frac{|x^{\prime}|}{1696\ \text{pc}}\right)^{C_{\perp}}+\left(\frac{|y^{\prime}|}{642.6\ \text{pc}}\right)^{C_{\perp}}$ (7)
$R_{S}^{C_{\parallel}}=R_{\perp}^{C_{\parallel}}+\left(\frac{|z^{\prime}|}{442.5\ \text{pc}}\right)^{C_{\parallel}}$ (8)
where $C_{\parallel}=3.501$ and $C_{\perp}=1.574$. The coordinates $x^{\prime},\ y^{\prime}$ and $z^{\prime}$ are Cartesian coordinates in the boxy bulge frame, which is rotated $13.79\degree$ around the $z$-axis and then $0.023\degree$ around the new $y$-axis, relative to the frame in which $x_{\odot}=-8300$ pc and $y_{\odot}=z_{\odot}=0$. See Fig. 1, right, for a plot of the boxy bulge distribution.
The second, spherical population of MSPs consists of $\rho_{\text{disk}}$ from Eq. (5) and a single, spherically symmetric density distribution, such that $\rho_{\text{spherical pop.}}=\rho_{\text{disk}}+\rho_{\text{sph}}$, given by:
@@
Sampling a 2D function, $f(x,y)$, requires splitting the function into two conditional functions, $f(x|y)$ and $f(y|x)$, as the MCMC method discussed above are 1-dimensional. For piece-wise functions f(x,y), the integral bounds for $f_{x}(x)$ and $f_{y}(y)$ are given by the piece-wise bounds.
@@ III.2 Sampling the Millisecond Pulsar Population
with z, r, $z_{0},\sigma_{R}$ defined as in Eq. (5). A comparison between the generated sample, plotted as a 2D histogram, and Eq. (5), plotted as a contour map, can be seen in Fig. 3. To produce both boxy and the spherical bulge populations, Eq.  (6) and Eq. (9) were also sampled via MCMC methods, and was confirmed to match with the density distribution pictured in Fig.  1 and Fig. 2.
@@ IV Ellipticity and Moment of Inertia Distribution
The ellipticity distribution was then folded with the morphological models described above, resulting in the projected pulsar ellipticity distribution shown in Fig. 5 for a randomly drawn sample of 40,000 MSPs for the Boxy (left) and Spherical (right) population models. The boxy distribution includes the Galactic disk component while the spherical distribution does not, resulting in the significant difference in morphology that can be observed. This is to better reflect the ambiguity of the GCE morphology as derived from the specific model of the gamma-ray backgrounds. As mentioned in Sec. I, due to uncertainties in the current interstellar gas maps and the background models used, a full determination of the morphology of the GCE has yet to be completely understood. Also note that the resultant ellipticity distribution is within the minimum ellipticity ranges described by Woan et al. Woan et al. (2018) and Chen et al. Chen (2020) of $\epsilon\approx 10^{-9}$, and agrees with the values obtained by Miller et al. Miller and Zhao (2023), as well as within the recent constraints given by Holst et al. Holst and Hooper (2024).
Figure 5: Locations and ellipticities of sampled MSPs for the boxy (left) and spherical (right) population models. The disk MSPs extend from -90 deg. to 90 deg. in the boxy population; the disk component is missing from the population model on the right, aimed at reproducing the possible GCE morphology.
@@ VI Analysis and Possibility of Detection
We note that if the GCE is produced by MSPs, a significant fraction thereof would exist within binary systems, as discussed in Sec. I, and therefore our assumptions that our population of MSPs are isolated or occur with no orbital eccentricity is a significant simplification. Depending on the orbital eccentricity of these binary systems, the MSP’s companion star could exert a significant enough tidal force to cause further deformations to the MSP, in addition to the magnetic field deformations already discussed. These could, in turn, significantly modify the accompanying GW amplitudes and frequencies, producing different results than discussed above. This may also necessitate a comparison to different gravitational wave detector sensitivities, as the detectors’ sensitivities in Fig. 10 and 11 correspond to the case of continuous gravitational wave sources from isolated objects, not objects in binary systems.
@@ VII Conclusion
In the future this analysis could be further strengthened by exploring the impact of a time-dependent ellipticity distribution and a frequency-dependent moment of inertia distribution on the strength and detectability of the GW signal. Additionally, it could be that the Galactic center excess is caused by a combination of MSPs and other, more exotic sources such as annihilating dark matter, or other unresolved astrophysical components, so it would be illuminating to incorporate such scenarios into the current model and distribution of MSPs and to evaluate the resulting GW signal. It should also be noted that Holst et al. Holst and Hooper (2024) suggests a range of possible MSP population sizes from 36,000-47,000 which, along with the unknown MSP luminosity function, suggests that the relation between the GCE and MSP population sizes is relatively poorly constrained. Therefore, further studies into how different population sizes affects detectability would be of great benefit. Finally, an investigation into the effects of binary systems and tidal deformations of MSPs on the resultant gravitational wave signal would give further useful insight into the possibility of using GWs to constrain the cause of the GCE.
Acknowledgments